OpenStudy (anonymous):

\[6 \over 5-\sqrt{3}\] \[6 \over 5-\sqrt{3}\] @Mathematics

OpenStudy (anonymous):

conjugate

OpenStudy (anonymous):

Multiply (5+sqrt(3)) at both denominator and numerator

jhonyy9 (jhonyy9):

30-3sqrt3 ---------- = (30-3sqrt3)/2 5-3

jhonyy9 (jhonyy9):

sorry (30-6sqrt3)/2

jhonyy9 (jhonyy9):

than is 2(15-3sqrt3)/2 = 15-3sqrt3

jhonyy9 (jhonyy9):

yes multiply numerator and denominator by 5+sqrt3

jhonyy9 (jhonyy9):

than will get (30+6sqrt3)/(5-3) = (30+6sqrt3)/2 =2(15+3sqrt3)/2 = 15+3sqrt3